OpenStudy (anonymous):

solve for X 5x^2+10x-495=0

OpenStudy (anonymous):

Start by factoring the expression: \[5(x ^{2}+2x-99)=0\] Then divide out the five to get\[(x ^{2}+2x-99)=0\] then you can factor again to get \[(x+11)(x-9)=0\] you then know that x+11=0 and that x-9 = 0 so you can get your solutions to x = -11 and x=9